Problem 16.14

A charge of 2.65 mC is placed at each corner of a square 0.170 m on a side.

Part A

Determine the magnitude of the force on each charge.

Thank you
Answer #1

45 4 S na 32 3

here , charge one each square ,

q = 2.65 mC

length of side , a = 0.170 m

as the magnitude of F1 = F2 = F

the net force is Fnet = 2 * F* cos(45) + F3

Fnet = 2 * cos(45) 9*10^9 * (2.65 *10^-3)^2/.170^2 + 9*10^9 * (2.65 *10^-3)^2/(0.170 * sqrt(2))^2

solving for Fnet

Fnet = 4.19 *10^6 N

the magntitude of force on each charge is 4.19 *10^6 N